8. The economy is experiencing 50 percent inflation per month. If a donut costs $1.50 today then how much it will cost next month. (Show calculations) (3 points)
Answer #1

Current price, P0 = $ 1.50

Rate of Inflation = 50% per month

Due to inflation the price after a month will be equal to

Price (after a month) = P0(1+Inflation)

Price = 1.5*(1+0.5) = 1.5 *1.5 = $ 2.25

After a month the donut will cost = $ 2.25

Please contact if having any query thank you.